OpenStudy (anonymous):

lim as x->0 (sin(7x))/x

OpenStudy (anonymous):

you need to remember that sinx/x is a squeeze theorem limit. does that help?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ \sin 7 x }{ 7 x }\times7=1\times7=7\]

OpenStudy (anonymous):

seems legit.

OpenStudy (anonymous):

but i think the op would get more out of it if he didnt get a complete solution for the price of no work.

OpenStudy (anonymous):

I don't get how squeeze theorem works…@lonnie455rich @surjithayer

OpenStudy (anonymous):

@lonnie455rich

OpenStudy (anonymous):

you take your trig function. and you set it in between its range. then perform the rest of the mathematical operations on the inequality. take the limit of the 2 outer terms in the inequality and if theyre equal then the limit is equal to the other 2 limits. I know im horrible at explaining. but its hard on a keyboard. ill try a more general way to explain it

OpenStudy (anonymous):

Ok, if you have some function. any function. and you say that the value is in between 2 functions. and the limit value of the outside functions are equal. then the inside function value limit must be at that point. Im sure a math major would shoot me for my wording. But thats the way i understand it. if you have \[1\le x \le1 \] what can you tell me about x?